I had the problem x-3y-8 and i found the slope first which is 1/3, but i have to use the negative recipricol of that so i used -3,then i had to find the point slope formula of that, i got y+4=-3(x-4) but now its asking me for the general form of the equation and idk how to do that!
Answer by jim_thompson5910(35256)

If you want it in slope intercept form, then


y+4=-3(x-4)

y+4=-3(x)-3(-4)

y+4=-3x+12

y=-3x+12-4

y=-3x+8

So the answer is y=-3x+8